OpenStudy (antonio_xx2):

Dave owns 15 shares of ABC mining stock. on monday the value of each shares rose $2 but on tuesday the value fell $5. what is the change in the value of dave's shares?

OpenStudy (perl):

2 - 5 = -3 dollars per share

OpenStudy (perl):

you could also solve it by total value of shares. the shares rose 15*2 total value = 30 dollars then it sank 5 dollars per share 15*(-5) = -75 30 - 75 = -45 total loss. divide that by total shares, that is 3 dollars loss per share

OpenStudy (antonio_xx2):

ok